OpenStudy (anonymous):

if the measures of the two angles have a sum of 90 degrees, then the angles are complimentary M Angle A + M Angle B = 90

OpenStudy (anonymous):

m=-a

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

m*a+m*b=90 degrees m is common let it go left is a+b=90 degrees so, 90/2=45 degrees by this we can say that => m=-a